OpenStudy (anonymous):

Simplify - 2\sqrt[3]{{11}} - \left( { - 8\sqrt[4]{3}} \right) - 8\sqrt[4]{3} + 15\sqrt[3]{{11}}. Write your answer in simplest radical form.

OpenStudy (unklerhaukus):

\[- 2\sqrt[3]{{11}} - \left( { - 8\sqrt[4]{3}} \right) - 8\sqrt[4]{3} + 15\sqrt[3]{{11}}.\]

OpenStudy (anonymous):

do u know how to simplify this ?

OpenStudy (unklerhaukus):

group the like terms \[- 2\sqrt[3]{{11}} - \left( { - 8\sqrt[4]{3}} \right) - 8\sqrt[4]{3} + 15\sqrt[3]{{11}}\\ =(-2+15)\sqrt[3]{11}+\big(-(-8)-8\big)\sqrt[4]{3}\\=\] then simplify the terms in brackets

OpenStudy (anonymous):

what did you end up

OpenStudy (unklerhaukus):

you tell me

OpenStudy (anonymous):

Im not smart.

OpenStudy (unklerhaukus):

(−2+15) = ?

OpenStudy (anonymous):

13

OpenStudy (unklerhaukus):

good , what about (−(−8)−8)

OpenStudy (unklerhaukus):

if it looks a bit confusing try to simplify this bit first -(-8)= ?

OpenStudy (anonymous):

-8?

OpenStudy (anonymous):

or just 8

OpenStudy (unklerhaukus):

-(-8)=+8

OpenStudy (unklerhaukus):

so what is (−(−8)−8) = ?

OpenStudy (anonymous):

64..?

OpenStudy (unklerhaukus):

we just said −(−8)=8 so (−(−8)−8) = (+8−8) = (8−8) = ?

OpenStudy (anonymous):

so i solve this ?

OpenStudy (anonymous):

u there

OpenStudy (unklerhaukus):

i am here,

OpenStudy (unklerhaukus):

what is eight minus eight?

OpenStudy (anonymous):

0

OpenStudy (unklerhaukus):

so , can you work out the final answer now/

OpenStudy (anonymous):

could u tell me the answer and explain the steps? that helps me a lot better

OpenStudy (unklerhaukus):

if you read all of the above posts slowly, you will be able to answer the question your self

OpenStudy (anonymous):

ok what about the 3 and 11 ? and 4 and 3

OpenStudy (anonymous):

???

OpenStudy (unklerhaukus):

those bits cannot be simplified , they will be in your final answer

OpenStudy (anonymous):

so, 13+0 then you sqaure the other numbers

OpenStudy (anonymous):

after u simplify

OpenStudy (unklerhaukus):

\[- 2\sqrt[3]{{11}} - \left( { - 8\sqrt[4]{3}} \right) - 8\sqrt[4]{3} + 15\sqrt[3]{{11}}\\ =(-2+15)\sqrt[3]{11}+\big(-(-8)-8\big)\sqrt[4]{3}\\=(13)\sqrt[3]{11}+\big(0\big)\sqrt[4]{3}\\=\] [dont square anything]

OpenStudy (anonymous):

dont you do 11*11*11

OpenStudy (anonymous):

next